# On biquandle-based invariant of immersed surface-links and on Yoshikawa oriented fifth move

### v2

@@ Abstract.
We resolve an open problem showing that the Yoshikawa’s fifth oriented move in his list cannot be reproduced by any finite sequence of the other nine moves and planar isotopies. Our proof introduces a link-type semi-invariant that remains unchanged under all moves except the fifth; contrasting values on two equivalent diagrams force the move’s independence. Second, we extend the algebraic toolkit for immersed surface-links. After revisiting the banded-unlink description of immersed surfaces and the twelve local moves that relate their diagrams, we develop a coloring theory based on biquandles. By assigning elements of a biquandle to diagram arcs according to local rules, we obtain a counting invariant of immersed surfaces up to isotopy.
@@
triple-crossing diagram
@@ 1. Introduction
The paper tackles two central problems in the diagrammatic study of oriented surfaces immersed in $\mathbb{R}^{4}$.
First, it settles an open question about Yoshikawa’s moves, the ten local moves that generate all diagrammatic isotopies of oriented surface-links. Previous work had proved the independence of every move except the oriented fifth move. We construct a semi-invariant derived from admissible marked-graph diagrams whose value is unchanged by every move in the generating set of other moves. Exhibiting two diagrams of the same surface-link with different values.
Second, the paper enriches the algebraic toolkit for immersed (i.e. generically self-intersecting) surface-links. We revisit the singular marked-graph (or “banded unlink”) presentation, extending the catalog of oriented planar moves so that any two diagrams of equivalent immersed surfaces differ by the twelve types of moves. We equip an immersed surface-link with a biquandle structure. Each semi-arc of a singular marked diagram is colored by elements of a biquandle according to local rules that respect crossings, markers and singular points. Because the biquandle axioms are tuned to the $\Gamma$-moves, the number of colorings is an invariant of the underlying surface. The authors compute this new biquandle coloring invariant for an illustrative example, obtaining two admissible colorings with a four-element biquandle.
Taken together, the independence proof sharpens our understanding of the move calculus, while the coloring invariant supplies fresh algebraic data for distinguishing immersed surfaces, advancing both the combinatorial and algebraic sides of four-dimensional knot theory.
This paper is organized as follows. In Section 2, we will prove an open problem of the independence of Yoshikawa oriented fifth move $\Gamma_{5}$ from the other generating set of oriented moves $\Gamma_{1},\ldots,\Gamma_{8},\Gamma_{4}^{\prime},\Gamma_{6}^{\prime}$ connecting diagrams of isotopic oriented surface-links.
In Section 3 we review the singular marked graph diagram method of presenting immersed surface-links and present the oriented planar moves between these diagrams.
In Section 4 we present the biquandle structure of an immersed surface-link with a method to obtain a biquandle coloring from a singular marked graph diagram and present a new invariant for oriented immersed surface-links. We show an example of the calculation.
@@ Acknowledgements
”This research was funded in whole or in part by NCN 2023/07/X/ST1/00157. For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.”
@@ 2. Independence of the Yoshikawa oriented fifth move
An embedding (or its image when no confusion arises) of a closed (i.e. compact, without boundary) surface $F$ into the Euclidean $\mathbb{R}^{4}$ (or into the $\mathbb{S}^{4}=\mathbb{R}^{4}\cup\{\infty\}$) is called a surface-link (or surface-knot if it is connected).
Two surface-links are equivalent if there exists an orientation preserving homeomorphism of the four-space $\mathbb{R}^{4}$ to itself (or equivalently auto-homeomorphism of the four-sphere $\mathbb{S}^{4}$), mapping one of those surfaces onto the other. See [Kam17] for an introductory material on surface-links.
@@
It is known that the set of ten types of moves $\mathcal{G}=\{\Gamma_{1},\ldots,\Gamma_{8},\Gamma_{4}^{\prime},\Gamma_{6}^{\prime}\}$ (presented in Fig. 1 and  2), called oriented Yoshikawa moves ([Yos94]), is a generating set of moves that relates two marked graph diagrams (modulo a planar isotopy) presenting equivalent oriented surface-links. In ([JKL13], [JKL15]) it is shown that any Yoshikawa move from the set $\mathcal{G}\backslash\{\Gamma_{5},\Gamma_{8}\}$ is independent of the other nine types. It is an open problem (see [JKL15], [Oht16]) whether the move $\Gamma_{5}$ is independent of the other Yoshikawa moves from the set $\mathcal{G}$.
@@
The Yoshikawa move $\Gamma_{5}$ cannot be realized by a finite sequence of Yoshikawa moves of the other nine types from the set $\mathcal{G}$, and planar isotopy.
@@
The diagrams $D_{1}$ and $D_{2}$ represent equivalent surface-links, with the desired property explained above, are illustrated in Figure 4, where we have that $L_{*}(D_{1})$ is ambient isotopy types of the $6_{1}$ knot (i.e. a nontrivial classical knot with the Alexander polynomial $2-5t+2t^{2}$), on the other hand, $L_{*}(D_{2})$ is ambient isotopy type of the trivial classical link.
@@ 3. Singular banded unlinks
When we consider the case of immersions $\mathbb{S}^{1}\to\mathbb{R}^{2}$ we have the ”classical” case, which includes, for example, planar projections of knots and links embedded in $\mathbb{R}^{3}$, loosing the (codimension two) knotting phenomena. The next-dimension case is $f:X^{2}\to\mathbb{R}^{4}$ when we have both the finite point generic intersections and nontrivial generic embeddings (a known generalization of the classical knot theory).
An immersion in this dimensions (or its image when no confusion arises) of a closed (i.e. compact, without boundary) surface $F$ into the Euclidean $\mathbb{R}^{4}$ (or into the $\mathbb{S}^{4}=\mathbb{R}^{4}\cup\{\infty\}$) is called a immersed surface-link (or immersed surface-knot if it is connected).
Two immersed surface-links are equivalent if there exists an orientation preserving homeomorphism of the four-space $\mathbb{R}^{4}$ to itself (or equivalently auto-homeomorphism of the four-sphere $\mathbb{S}^{4}$), mapping one of those surfaces onto the other. Fix an immersed surface-link $F$ embedded in a manifold $\mathbb{S}^{4}$. For an open neighborhood, denoted $N(F)$, the exterior of $F$ is $E(F):=\mathbb{S}^{4}\backslash N(F)$.
@@ 4. Biquandle structure
A generalization of quandles (called biquandles) was introduced in [KauRad03]. A biquandle is an algebraic structure with two binary operations satisfying certain conditions which can be presented by semi-arcs of links (or semi-sheets of surface-links) as its generators modulo oriented Reidemeister moves (or Roseman moves). In [CES04], J. S. Carter, M. Elhamdadi, and M. Saito introduced and used cocycles to define invariants via colorings of link diagrams by biquandles and a state-sum formulation.
@@
Let $X$ and $Y$ be biquandles. A function $f:X\rightarrow Y$ is called a biquandle homomorphism if $f({x}\,\underline{\triangleright}\,{y})={f(x)}\,\underline{\triangleright}\,{f(y)}$ and $f({x}\,\overline{\triangleright}\,{y})={f(x)}\,\overline{\triangleright}\,{f(y)}$ for any $x,y\in X.$ We denote the set of all biquandle homomorphisms from $X$ to $Y$ by ${\rm Hom}(X,Y)$. A bijective biquandle homomorphism is called a biquandle isomorphism. Two biquandles $X$ and $Y$ are said to be isomorphic if there is a biquandle isomorphism $f:X\to Y$.
@@
The biquandle axioms are chosen such that given a biquandle coloring of one side of any $\Gamma$–type move, there is a unique biquandle coloring of the other side of the move with the condition that colors agree on the boundary arcs that leaves the disc where the move is performed.
@@
The proof for the moves $\Gamma_{1},\ldots,\Gamma_{3}$ can be found in the existing literature (see [KNS16]), the axioms for the biquandle were motivated to satisfy those Reidemeister moves, for moves $\Gamma_{4},\ldots,\Gamma_{8}$ see for example [JouNel20]. We checked, by standard hand calculations, the proof for the remaining moves, using the diagrams shown in Figures 12, 911.
@@
We can enumerate that #$\#Col_{X_{T}}(D_{T})=5$, two of the admissible coloring is shown in Figure 13, the other two admissible colorings are obtained from the previous by exchanging colors $1$ with $3$, and the fifth one is monochromatic, colored by element $2$.

### v3

# On biquandle-based invariant of immersed surface-links, Yoshikawa oriented fifth move, and ribbon 2-knots

@@ Abstract.
We resolve an open problem by showing that the Yoshikawa’s fifth oriented move in his list cannot be reproduced by any finite sequence of the other nine moves and planar isotopies. Our proof introduces a link-type semi-invariant that remains unchanged under all moves except the fifth, highlighting its necessity in generating the full move set. Second, we extend the algebraic toolkit for immersed surface-links. After revisiting the banded-unlink description of immersed surfaces and the twelve local moves that relate their diagrams, we develop a biquandle-based coloring theory. By assigning elements of a biquandle to diagram arcs according to local rules, we obtain a counting invariant of immersed surfaces up to isotopy. Third, we show that there are infinitely many pairs of ribbon $2$-knots with isomorphic groups but different knot quandles.
@@
biquandle invariant, immersed surface-links, Yoshikawa moves, ribbon 2-knots
@@ 1. Introduction
The paper settles an open question about Yoshikawa’s moves, the ten local moves that generate all diagrammatic isotopies of oriented surface-links. Previous work had proved the independence of other moves, not the oriented fifth move. We construct a semi-invariant derived from admissible marked-graph diagrams whose value is unchanged by every move in the generating set of other moves. Exhibiting two diagrams of the same surface-link that yield different values of the semi-invariant.
Second, we enrich the algebraic framework for immersed (i.e., possibly self-intersecting) surface-links. We revisit the singular marked-graph (or ”banded unlink”) presentation, extending the catalog of oriented planar moves so that any two diagrams of equivalent immersed surfaces differ by the twelve types of moves. We equip an immersed surface-link with a biquandle structure. Each semi-arc of a singular marked diagram is colored by elements of a biquandle according to local rules that respect crossings, markers and singular points. Because the biquandle axioms correspond to the $\Gamma$-moves, the number of colorings is an invariant of the underlying surface type. The authors compute this new biquandle coloring invariant for an illustrative example, obtaining admissible colorings with a three-element biquandle.
Third, we prove that there are infinitely many pairs of inequivalent ribbon $2$-knots that share the same fundamental group of the exterior but differ in the fundamental quandle.
This paper is organized as follows. In Section 2, we will prove an open problem of the independence of Yoshikawa oriented fifth move $\Gamma_{5}$ from the other generating set of oriented moves $\Gamma_{1},\ldots,\Gamma_{8},\Gamma_{4}^{\prime},\Gamma_{6}^{\prime}$ connecting diagrams of isotopic oriented surface-links. In Section 3, we review the singular marked graph diagram method of presenting immersed surface-links and present the oriented planar moves between these diagrams. In Section 4, we present the biquandle structure of an immersed surface-link with a method to obtain a biquandle coloring from a singular marked graph diagram and present a new invariant for oriented immersed surface-links. We show an example of the calculation. In Section 5, we show that there are infinitely many pairs of ribbon $2$-knots with isomorphic groups but different knot quandles.
@@ Acknowledgements
This research was funded in whole or in part by NCN 2023/07/X/ST1/00157.
@@ 2. Independence of the Yoshikawa oriented fifth move
An embedding (or its image when no confusion arises) of a closed (i.e., compact, without boundary) surface $F$ into the Euclidean $\mathbb{R}^{4}$ (or into the $\mathbb{S}^{4}=\mathbb{R}^{4}\cup\{\infty\}$) is called a surface-link (or surface-knot if it is connected).
Two surface-links are equivalent if there exists an orientation-preserving homeomorphism of the four-space $\mathbb{R}^{4}$ to itself (or equivalently an auto-homeomorphism of the four-sphere $\mathbb{S}^{4}$), mapping one of those surfaces onto the other. See [Kam17] for introductory material on surface-links.
@@
It is known that the set of ten types of moves $\mathcal{G}=\{\Gamma_{1},\ldots,\Gamma_{8},\Gamma_{4}^{\prime},\Gamma_{6}^{\prime}\}$ (presented in Fig. 1 and  2), called oriented Yoshikawa moves ([Yos94]), is a generating set of moves that relates two marked graph diagrams (modulo a planar isotopy) presenting equivalent oriented surface-links. In [JKL13], [JKL15] it is shown that any Yoshikawa move from the set $\mathcal{G}\backslash\{\Gamma_{5},\Gamma_{8}\}$ is independent of the other nine types. It is an open problem (see [JKL15], [Oht16]) whether the move $\Gamma_{5}$ is independent of the other Yoshikawa moves from the set $\mathcal{G}$.
@@
The Yoshikawa move $\Gamma_{5}$ cannot be realized by a finite sequence of Yoshikawa moves of the other nine types from the set $\mathcal{G}$, and a planar isotopy.
@@
The diagrams $D_{1}$ and $D_{2}$ represent equivalent surface-links, with the desired property explained above, are illustrated in Figure 4, where we have that $L_{*}(D_{1})$ is of the ambient isotopy type of the $6_{1}$ knot (i.e. a nontrivial classical knot with the Alexander polynomial $2-5t+2t^{2}$), on the other hand, $L_{*}(D_{2})$ is ambient isotopy type of the trivial classical link.
@@ 3. Singular banded unlinks
When we consider the case of immersions $\mathbb{S}^{1}\to\mathbb{R}^{2}$ we have the ”classical” case, which includes, for example, planar projections of knots and links embedded in $\mathbb{R}^{3}$, losing the (codimension two) knotting phenomena. The next-dimension case is $f:X^{2}\to\mathbb{R}^{4}$ when we have both the finite point generic intersections and nontrivial generic embeddings (a known generalization of the classical knot theory).
An immersion in this dimensions (or its image when no confusion arises) of a closed (i.e. compact, without boundary) surface $F$ into the Euclidean $\mathbb{R}^{4}$ (or into the $\mathbb{S}^{4}=\mathbb{R}^{4}\cup\{\infty\}$) is called an immersed surface-link (or immersed surface-knot if it is connected).
Two immersed surface-links are equivalent if there exists an orientation-preserving homeomorphism of the four-space $\mathbb{R}^{4}$ to itself (or equivalently an auto-homeomorphism of the four-sphere $\mathbb{S}^{4}$), mapping one of those surfaces onto the other. Fix an immersed surface-link $F$ embedded in a manifold $\mathbb{S}^{4}$. For an open neighborhood, denoted $N(F)$, the exterior of $F$ is $E(F):=\mathbb{S}^{4}\backslash N(F)$.
@@ 4. Biquandle structure
A generalization of quandles (called biquandles) was introduced in [KR03]. A biquandle is an algebraic structure with two binary operations satisfying certain conditions which can be presented by semi-arcs of links (or semi-sheets of surface-links) as its generators modulo oriented Reidemeister moves (or Roseman moves). In [CES04], J. S. Carter, M. Elhamdadi, and M. Saito introduced and used cocycles to define invariants via colorings of link diagrams by biquandles and a state-sum formulation.
@@
Let $X$ and $Y$ be biquandles. A function $f:X\rightarrow Y$ is called a biquandle homomorphism if $f({x}\,\underline{\triangleright}\,{y})={f(x)}\,\underline{\triangleright}\,{f(y)}$ and $f({x}\,\overline{\triangleright}\,{y})={f(x)}\,\overline{\triangleright}\,{f(y)}$ for any $x,y\in X.$ We denote the set of all biquandle isomorphic from $X$ to $Y$ by ${\rm Hom}(X,Y)$. A bijective biquandle homomorphism is called a biquandle isomorphism. Two biquandles $X$ and $Y$ are said to be isomorphic if there is a biquandle isomorphism $f:X\to Y$.
@@
The biquandle axioms are chosen such that given a biquandle coloring of one side of any $\Gamma$–type move, there is a unique biquandle coloring of the other side of the move with the condition that colors agree on the boundary arcs that leave the disc where the move is performed.
@@
The proof for the moves $\Gamma_{1},\ldots,\Gamma_{3}$ can be found in the existing literature (see [KNS16]), the axioms for the biquandle were motivated to satisfy those Reidemeister moves, for moves $\Gamma_{4},\ldots,\Gamma_{8}$ see for example [JN20]. We checked, by standard hand calculations, the proof for the remaining moves, using the diagrams shown in Figures 12, 911.
@@
We find that #$\#Col_{X_{T}}(D_{T})=5$, two of the admissible coloring is shown in Figure 13, the other two admissible colorings are obtained from the previous by exchanging colors $1$ with $3$, and the fifth one is monochromatic, colored by element $2$.
@@
5. Ribbon $2$-knots and their fundamental quandles
Let us consider an oriented surface-knot $F$ in $\mathbb{S}^{4}$. From Tubular Neighborhood Theorem $F$ admits a tubular neighborhood, denoted $N(F)$, i.e., an image of a smooth embedding $N:\mathbb{D}^{2}\times F$ such that for any $p\in F$ we have $N(0,p)=p$. The exterior of $F$ is $E(F):=\mathbb{S}^{4}\backslash int(N(F))$. If two surface-knots are equivalent, then their exteriors are diffeomorphic. The group of a surface-link $F$ denoted $\mathcal{G}(F)$ is $\pi_{1}(E(F))$, the group is finitely presented.
The other invariant of surface-knot $F$ is its fundamental quandle $\mathcal{Q}(F)$, associated group of the knot quandle is group isomorphic to the knot group (see [Kam17] for an introduction). Every $2$-knot that has a ($\mathbb{R}^{3}$) diagram without triple points is ribbon [Yaj64]. Two ribbon $2$-knots or ribbon torus-knots with isomorphic fundamental quandles have isomorphic fundamental biquandles [Ash14]. Therefore, we focus in this section on fundamental quandles, for this family of $2$-knots. It is known that there exist arbitrarily many inequivalent surface-knots with the same knot fundamental quandle [Tan07].
Recently, in [TT25] it is shown that there exist infinitely many pairs of $2$-knots not distinguishable by a knot group but tell apart by their quandles. However, none of those pairs were both ribbon $2$-knots, at least one of them was non-ribbon twist-spun knot. It is natural to ask a question if there is a pair of ribbon $2$-knots with isomorphic groups that have different knot quandles. We obtain the following.
Theorem 5.1.
There are infinitely many pairs of ribbon $2$-knots with mutually isomorphic fundamental groups of their exteriors but mutually non-isomorphic knot fundamental quandles.
Proof.
Suciu [Suc85] constructed an infinite family of mutually non-ambient-isotopic $2$-knots (with non-isomorphic $\pi_{2}$ as $\mathbb{Z}\pi_{1}$ modules), denoted $R_{k}$, for any integer $k>0$. We show that this family of surface-knots has the desired property in the Theorem. Each $\mathcal{G}(R_{k})$ is isomorphic to the trefoil classical knot group $t,x\;|\;txt=xtx\right>$.
First, let us derive the presentation for the fundamental biquandle $\mathcal{Q}(R_{k})$. The presentation for the fundamental biquandle can be calculated directly from the marked graph diagram for surface-knot [Ash12]. It agrees with the method described earlier in this paper. Specifically, we consider relations as in Figure 12, but without singular points and with the trivial one of the operations, namely $x\,\overline{\triangleright}\,y=x$ for all $x,y$.
We present a marked graph diagram for $R_{k}$ in Figure 14, which is the modified, by the relation in Figure 6, version of the banded unlink diagram for $R_{k}$ presented in [Rup22, p. 55]. In the Figure, the polygonal green region has to appear $k-1$ times to obtain $R_{k}$. From the arc relations, we see that if we put arc generators $c,d,t$, then we end up with two relations, marked by the red rectangles.
{lpic}
[]sc2_5_quandle(14.5cm)
Figure 14. A marked graph diagrams for $R_{k}$.
Therefore, for any integer $k>0$:
$\mathcal{Q}(R_{k})=\left<c,d,t\;|\;(d*t)*d=(c*t)*c\;,\;(d*t)*c=(\ldots(t*d)\overline{*}c)*d\ldots)\overline{*}c)*d\right>,$
where $\overline{*}$ is the dual quandle operation to *, and in the second dots $\ldots$ symbols ”)¯$)\overline{*}c)*d$” appear $k-2$ times.
It is of no use here to consider the quandle cocycle invariant (see [Kam17] for the definition) to distinguish any pair of them. This is because the quandle cocycle invariant can be calculated from the neighborhoods of the triple points in the projection, and each ribbon $2$-knot has a projection without any triple point (as we stated earlier).
We know that $\mathcal{G}(R_{i})\cong\mathcal{G}(R_{j})$ for any $i\not=j$. Note that it is also of no use here to consider the associated group $As(Q)=\left<x\in Q\;|\;x*y=y^{-1}xy\;(\text{for }x,y\in Q)\right>$, because $As(\mathcal{Q}(R_{i}))\cong\mathcal{G}(R_{i})\cong\mathcal{G}(R_{j})\cong As(\mathcal{Q}(R_{i}))$ for any $i\not=j$. Our idea to show that $\mathcal{Q}(R_{i})\not\cong\mathcal{Q}(R_{j})$ for any $i\not=j$ is to use the core groups of their fundamental quandles. The core group of the quandle $Q$ is a group with presentation $Core(Q)=\left<x\in Q\;|\;x*y=yx^{-1}y\;(\text{for }x,y\in Q)\right>$.
We can straightforward calculate their presentations (for any integer $k>0$) with substitutions ($t=t^{-1}$, $w=dc^{-1}$):
$Core(\mathcal{Q}(R_{k}))=\left<c,t,w\;|\;wctwctw=ctct\;;\;ctwct=w^{k+1}ctw^{k+1}\right>.$
To show that these groups are mutually non-isomorphic, we use pure group theory [MKS04, chapter 5], [Rob96, chapter 5.1].
Denote $A_{k}=Core(\mathcal{Q}(R_{k}))$ for short, we want to show that $A_{n}\not\cong A_{m}$ for any $n\not=m$. In the abelianization $A_{k}/[A_{k},A_{k}]\cong\mathbb{Z}\oplus\mathbb{Z}_{3}$ the two defining relations becomes $03[w]=$ 0, $[c]+[t]=(2k+1)[w]$. The order of $[c]+[t]$ in the torsion summand is $3/gcd(3,2k+1)$.
A direct Hall-Witt computation gives
$[w^{3},c]=[[w,c],w]^{3}\ [[w,c],w,c],\qquad[w^{3},t]=[[w,t],w]^{3}\ [[w,t],w,t],$
so $w^{3}$ is central whenever every weight-2 commutator with $w$ has exponent dividing $3$.
Therefore $w^{3}$ is central iff $k\equiv 1\pmod{3}$. So if $n\not\equiv m\pmod{3}$ then $A_{n}\not\cong A_{m}$.
Assume that $n\equiv m\pmod{3}$. Consider group lower-central series $\gamma_{1}(G)=G,\gamma_{s+1}(G)=[\gamma_{s}(G),G]$. Assume that (in the contrary) $A_{n}\cong A_{m}$, then because $\gamma_{s}(G)$ is characteristic we have $A_{n}/\gamma_{4}(A_{n})\cong A_{m}/\gamma_{4}(A_{m})$. With the standard Hall commutator calculus, only the triple commutator $[[c,t],w]$ survives.
[…]